OpenStudy (please.help.me):

Find the exact value of the trigonometric expression given that sin u =8/17 and cos v = −15/13 (Both u and v are in Quadrant II.) sin(u + v)

satellite73 (satellite73):

it is impossible for cosine to be equal to \(-\frac{15}{13}\) did you mean \(-\frac{12}{13}\)?

OpenStudy (please.help.me):

@satellite73 I copied it exactly how it was written.

OpenStudy (please.help.me):

@3mar This is the one that I am talking about

satellite73 (satellite73):

then it is a mistake cosine cannot be smaller than \(-1\)

satellite73 (satellite73):

you got a screen shot or something to look at ?

OpenStudy (please.help.me):

This is a screenshot

satellite73 (satellite73):

hmm entirely different question

OpenStudy (please.help.me):

@satellite73 I posted a screenshot above

satellite73 (satellite73):

that is something about tangent, not the question you posted above

OpenStudy (please.help.me):

Sorry I uploaded the wrong one, hold on

satellite73 (satellite73):

ok holding i bet $7 it is \(-\frac{12}{13}\)

OpenStudy (please.help.me):

@satellite73

satellite73 (satellite73):

ok l lose but it is \(-\frac{5}{13}\) not \(-\frac{15}{13}\) whew

satellite73 (satellite73):

now we can do it

satellite73 (satellite73):

\[\sin(u+b)=\sin(u)\cos(v)+\sin(v)\cos(u)\] you already know two out of the four numbers, you need the other two

OpenStudy (please.help.me):

You just wrote exactly the same numbers

satellite73 (satellite73):

in inglish minus five over thirteen, not minus fifteen over thirteen

OpenStudy (please.help.me):

Do you mean sin(u+V)=sin(u)cos(v)+sin(v)cos(u) ?

OpenStudy (3mar):

@satellite73 is right! It is not logic and not what we know!

satellite73 (satellite73):

yes you know two out of the four, you need the other two

OpenStudy (please.help.me):

Yes

satellite73 (satellite73):

\[\sin(u)=\frac{8}{17}\] what is \(\cos(u)\)?

satellite73 (satellite73):

or is that what you need help finding?

OpenStudy (please.help.me):

Yes, I am not really here how to get the cos

satellite73 (satellite73):

draw a right triangle is one easy way to do it

satellite73 (satellite73):



satellite73 (satellite73):

there is a picture of an angle whose sine is \(\frac{8}{17}\) what you need is the "adjacent" side, which you find via pythagoras

OpenStudy (please.help.me):

So 15 for the other side

satellite73 (satellite73):

yes, and therefor \[\cos(u)=-\frac{15}{17}\] negative because they told you you were in quadrant II

OpenStudy (please.help.me):

Why would it be negative though?

OpenStudy (please.help.me):

oh sorry just read it

satellite73 (satellite73):

lol

satellite73 (satellite73):

now we repeat the process to find \(\sin(v)\) given that \(\cos(v)=-\frac{5}{13}\)

satellite73 (satellite73):



OpenStudy (please.help.me):

12/5 ?

satellite73 (satellite73):

no that is tangent the 12 is right though

satellite73 (satellite73):

opposite over hypotenuse

OpenStudy (please.help.me):

Wait 5/13 sorry

satellite73 (satellite73):

no, that was what you were given, cosine you want sine

OpenStudy (please.help.me):

12/13

satellite73 (satellite73):

yes

satellite73 (satellite73):

now plug the numbers in and do the arithmetic, that is all that is left to do

satellite73 (satellite73):

\[\sin(u+v)=\sin(u)\cos(v)+\sin(v)\cos(u)\] you now know all four numbers on the right

OpenStudy (please.help.me):

Okay I am working on it

satellite73 (satellite73):

ok arithmetic is sometimes annoying

OpenStudy (please.help.me):

How do you simplify it though? Do you need common denominators?

satellite73 (satellite73):

if by "simplify" you mean find the number, yes, but you don't need to find a common denominator in this case because the denominators are equal

OpenStudy (please.help.me):

But it is split by a + sign so I can't just add the common terms together

satellite73 (satellite73):

the denominator of both the fraction on the right and the fraction on the left is \(13\times 17=221\)

satellite73 (satellite73):

what did you write down to compute?

OpenStudy (please.help.me):

\[\sin \frac{ 8 }{ 17 } \cos-\frac{ 5 }{ 13} + \sin \frac{ 12 }{ 13 } \cos-\frac{ 15 }{ 17 }\]

satellite73 (satellite73):

hmm i was afraid of that

satellite73 (satellite73):

it is not the sine OF \(\frac{8}{17}\) the sine IS \(\frac{8}{17}\)

satellite73 (satellite73):

\[ \frac{ 8 }{ 17 } \times -\frac{ 5 }{ 13} + \frac{ 12 }{ 13 } \times -\frac{ 15 }{ 17 }\]

satellite73 (satellite73):

those numbers ARE the sine and cosine of u and v you do not take the sine and cosine OF those numbers just compute with the numbers

OpenStudy (please.help.me):

What does compute mean? Sorry for asking

satellite73 (satellite73):

find the number? multiply and add i guess

satellite73 (satellite73):

perhaps what i really meant was "do the arithmetic"

OpenStudy (please.help.me):

\[-\frac{ 40 }{ 221 }+\] \[\frac{ 180 }{ 221}\]

OpenStudy (please.help.me):

-220/221 ?

satellite73 (satellite73):

looks right to me